# Auto-GDA: Automatic Domain Adaptation For Grounding Verification in Retrieval Augmented Generation

### v1

@@ 4.4 Automatic Selection of High-Quality Samples
Figure 4: Modeling the label correctness term in Eqn. 2 as function of $r$. When the estimated entailment certainty $r$ does not match the assigned hard label $\hat{y}$ this term takes high values discouraging selection.
@@
Theoretical Properties. Notably, Equation 2 can be derived from first principles as an enhanced distribution matching objective. By defining parametric distributions $p_{\mathcal{Q},{\bm{e}}}({\bm{c}},y)$ (representing the selected synthetic data for evidence ${\bm{e}}$) and $p_{\text{cov},{\bm{e}}}({\bm{c}},y)$ (representing the target distribution for ${\bm{e}}$ we aim to imitate) the objective corresponds to the divergence between these distributions plus the expected utility of the synthetic data. Formally,
@@
Optimizing the objective. Optimizing the objective for a subset $\mathcal{Q}$ of size $K$ with minimal loss can be done highly efficiently in three steps: (1) Computing each samples’ contribution to the sum in $\mathcal{L}_{tot}$, (2) ranking the samples by this contribution, and (3) greedily selecting the top-$K$ subset of samples with the lowest contributions. Pseudocode of our complete framework is provided in Algorithm 1 (Appendix).
@@ 5 Experimental Evaluation
Datasets. We evaluate our approach on three datasets for document-grounded summarization and question answering (QA). We select datasets which include documents, realistic LLM-generated long-form answers, and human labels that can be used for testing. The SummEdits (Laban et al., 2023) dataset contains GPT-3.5-generated and manual summaries of documents from different domains, e.g., judicial, sales emails, podcasts. We further use both the summary and the QA portion of the RAGTruth (Wu et al., 2023) dataset. The RAGTruth dataset contains summaries and answers to questions created by LLMs (GPT-3.5/4, Mistral, Llama2). Finally, we use the LFQA-Verification dataset (Chen et al., 2023a), which retrieved documents for questions from the “Explain me Like I am five”-dataset and generated corresponding long-form answers with GPT-3.5 and Alpaca. We selected the datasets to feature characteristics of realistic RAG systems including specific prompt templates (present in RAGTruth, LFQA-Verification) and various domains (present in all datasets, specifically in SummEdits). Details and links to the datasets can be found in Table 4.
@@
Baselines. We use state-of-the-art baselines: AlignScore (Zha et al., 2023) (RoBERTa-based with pre- and postprocessing), MiniCheck (Tang et al., 2024), and Vectara-2.1111https://docs.vectara.com/docs (both T5-based). As a teacher model to assign initial score, we use the best-performing model from the “complex” category, which allow easy access to uncertainty scores and have good performance. We employ optuna [^2] https://optuna.org/ to tune the remaining hyperparameters $\lambda_{u}^{\prime}$, $\lambda_{d}$, and the other teacher model used to estimate entailment probabilities for augmentations in Eqn. 1 performing 50 trials per dataset. Auto-GDA is run for two iterations on RAGTruth and one iteration on the other datasets, generating synthetic datasets between 1.3×1.3$\times$1.3 × and 2×2$\times$2 × the original dataset size. We found no improvements through further increasing dataset size. We also compare against several common UDA methods, including robustness-based approaches. Specifically, we implement DAPT (Gururangan et al., 2020), SiFT (He et al., 2020), and Deep-CORAL (Sun & Saenko, 2016) for further pretraining of the DeBERTa-V2 model. Additional implementation details for all methods can be found in Appendix C.
@@ 5.1 Synthetic Data for NLI Model Fine-Tuning
We present the main results obtained with Auto-GDA in Table 1. Our results show that Auto-GDA is highly effective and improves performance in ROC-AUC scores of all tested models on all our datasets. Additional metrics confirm our findings (balanced accuracy, f1-scores in Section D.1).
@@ 5.2 Ablation Investigations
Components of the algorithm. We add the components of our algorithm individually and show how they successively increase performance in Table 2. In all ablations we keep dataset size and other parameters constant. The biggest gain is achieved by fine-tuning on data created by few-shot prompting. We subsequently add data augmentations without applying our selection criterion, but instead selecting few-shot and augmented samples randomly. We observe that this decreases data quality, highlighting that data augmentation is only beneficial together with our filtering criterion. When we do so and apply data augmentation with our filtering step (corresponding to full Auto-GDA), this increases performance overall with one exception on the LFQA-Verification dataset (note however that performance here is already above the labeled data, so selection based on target data may draw the results toward the labeled data scores as well). As an upper baseline we are interested in the hypothetical performance reachable by fine-tuning on human-labeled samples and include it in Table 2. Considering the difference between the no fine-tuning models and the models fine-tuned on human-labeled data as the domain adaptation gap, expressing our results relative to these baselines indicates that we manage to close an impressive 96% of this gap.
@@
Selection from Several Augmentation Routines. We report the effect of single versus several augmentations in Figure 6(a) and Figure 6(b) in Section D.2, quantitatively and qualitatively demonstrating that Auto-GDA succeeds to the most promising samples from different augmentations.
@@ 5.3 Inference Efficiency
Linking to our motivational Figure 1, we study the efficiency of our models in Table 3. We compute NLI scores for 50 random samples from the respective datasets. We observe models in three categories: The most efficient models (Vectara to BART-large) have medium performance on the RAG datasets used in this work indicated by their ROC. On the other hand, models using sophisticated post-processing (AlignScore) perform better, but require about 2.5 times more inference time than our most successful DeBERTa model. Finally, LLMs via APIs require about 10-fold inference time, but result in highest performance. When we compare models trained with our approach, we observe LLM-level performance at about 10% of the inference time.
@@ 6 Disussion and Conclusion
In this work, we show that synthetic data can successfully tackle the domain generalization gap for NLI models. We present Auto-GDA, an automatic approach for synthetic sample generation and selection overcoming the need for tedious heuristic or manual filtering and augmentation selection. Our results show that we can obtain models that perform on par with most powerful LLMs while having around 90% less inference time using our method. Our findings further demonstrate the superiority of synthetic data generation over unsupervised domain adaptation (UDA) methods. Even with synthetic data obtained only through few-shot prompting, we can obtain better results than the UDA baselines (Table 1 vs. Table 2). We attribute this to the inherent limitations of UDA methods that usually operate in vectorial embedding spaces and fail to capture semantics of the target domain. By generating synthetic data, we can provide a more comprehensive and tailored representation, allowing for greater control over the desired features. Our results also confirm the common intuition that generalization is increasingly hard with smaller models (Bhargava et al., 2021). This highlights that domain adaptation is particularly important when low latency at inference time is required, whereas general purpose models can be preferable when quick inference is no hard requirement.
Limitations. In our study we assume that the distribution of evidence samples including the retrieved documents is readily available. In many real-world applications, this may not be the case. To address this, techniques like passage clustering and summarization, as explored in Sarthi et al. (2024), could be employed on the knowledge base to cover the diversity of evidence passages as a surrogate of this distribution. In addition, domain adaptation requires a model for each individual domain. Future work is required to further study if it is possible to adapt efficient models for multiple domains without performance degradation.
@@ B.1 Deriving our Objective from a Distribution Matching Criterion
Parametrizing densities. We need to insert a suitable and tractable parametrizations of $p_{\text{cov}}$ and $p_{\theta}$. We start by modeling their marginals. To model $p_{\text{cov}}$ we chose an efficiently tractable density $p_{\text{cov}}({\bm{x}})$ can be defined via the nearest target feature vector in $\text{claim}(\mathcal{D}_{t,{\bm{e}}})$ by333we now assume ${\bm{x}}\in\mathcal{X}$ to be in a metric space. This can be achieved using an encoder mapping the textual input to real vectors.
@@ C.1 Implementation Details for Augmentation Strategies
In this section we provide additional details regarding the data augmentation strategies that we deploy in this work.
@@ C.3 Auto-GDA Details and Hyperparameters
Base models. As a basis for finetuning we use huggingface checkpoints for DeBERTaV2666https://huggingface.co/tasksource/deberta-base-long-nli, BART-large [^7] https://huggingface.co/facebook/bart-large-mnli and FLAN-T5888https://huggingface.co/sjrhuschlee/flan-t5-base-mnli.
@@ C.5 Baseline NLI Models
For the complex NLI model baselines, we use Vectara HHEM-2.1999https://huggingface.co/vectara/hallucination_evaluation_model. The model cannot be easily fine-tuned because it uses custom code. Additionally we use Alignscore-base with the checkpoint found in this repository [^10] https://github.com/yuh-zha/AlignScore with the recommended “split” (pre- and postprocessing) nli_sp option. We neglect the larger version as its runtime was comparable to LLMs at a usually lower performance, making the smaller model a better trade-off. Finally we use the best-performing Minicheck flan-t5-large model by Tang et al. (2024) from the official huggingface page [^11] https://huggingface.co/lytang/MiniCheck-Flan-T5-Large.
@@ C.6 Robust Pre-training and Fine-tuning for Unsupervised Domain Adaptation
Implementation Details For our robust optimization experiments we used the DeBERTaV2 based NLI model, and limited maximum tokenization length at 1024 tokens across all benchmarks. For DAPT we extracted the DeBERTaV2 backbone and trained on the target domain for 1 full epoch, using 10% masking probability. At the fine-tuning stage of both methods we ran 1 full epoch on the MNLI dataset. At the masked pretraining and fine-tuning stages of the experiments, we used the AdamW optimizer with learning rate $10^{-5}$ and weight decay $10^{-3}$, enabling 100 warmup steps over the supervised fine-tuning. For SiFT and CORAL we set the coefficient of the respective regualarization terms to 0.5, after running hyperparameter optimization with a coarse grid search. Batch size used for covariance estimation in CORAL was set at 64. Each experiment was repeated over 3 random trials.
@@ D.3 Different teacher models
D.4 Simple Baselines
@@
Dataset RAGTruth LFQA-Verif. SummEdits Mean (Gap closed)
@@
Table 12: fine-tuning on validation set.
D.5 Effect of Dataset Size
@@
Figure 8: Testing different synthetic dataset sizes and learning rates for fine-tuning. While the original dataset size is at about 3 claim examples per evidence (dashed line) we test a wide range of dataset sizes ranging from 1/50 examples per evidence (100x smaller than original) to larger datasets with 30 claims per evidence, (10x larger than original). While larger We observe relatively stable optimum from about a 1/3 of original size to twice the original size using learning rates around $10^{-5}$.
D.6 Effect of Source Data on UDA effectivenss

### v2

@@ 4.4 Automatic Selection of High-Quality Samples
Figure 4: Modeling the label correctness term in Eqn. 2 as function of $r$. When the estimated entailment certainty $r$ does not match the assigned hard label $\hat{y}$ this term takes high values, discouraging selection.
@@
Theoretical Properties. Notably, Equation 2 can be derived from first principles as an enhanced distribution matching objective. By defining parametric distributions $p_{\mathcal{Q},{\bm{e}}}({\bm{c}},y)$ (representing the selected synthetic data for evidence ${\bm{e}}$) and $p_{\text{cov},{\bm{e}}}({\bm{c}},y)$ (representing the target distribution for ${\bm{e}}$ we aim to imitate) the objective corresponds to the divergence between these distributions minus the expected utility of the synthetic data. Formally,
@@
Optimizing the objective. Optimizing the objective for a subset $\mathcal{Q}_{\bm{e}}$ containing $K$ synthetic samples for evidence ${\bm{e}}$ with minimal loss can be done highly efficiently in three steps: (1) Computing each samples’ contribution to the sum in $\mathcal{L}_{tot}$, (2) ranking the samples by this contribution, and (3) greedily selecting the top-$K$ subset of samples with the lowest contributions. Pseudocode of our complete framework is provided in Algorithm 1 (Appendix).
@@ 5 Experimental Evaluation
Datasets. We evaluate our approach on three datasets for document-grounded summarization and question answering (QA). We select datasets that include documents, realistic LLM-generated long-form answers, and human labels that can be used for testing. The SummEdits dataset (Laban et al., 2023) contains GPT-3.5-generated and manual summaries of documents from different domains, e.g., judicial, sales emails, podcasts. We further use both the summary and the QA portion of the RAGTruth dataset (Niu et al., 2024). The RAGTruth dataset contains summaries and answers to questions created by LLMs (GPT-3.5/4, Mistral, Llama2). Finally, we use the LFQA-Verification dataset (Chen et al., 2023a), which retrieved documents for questions from the “Explain me Like I am five”-dataset and generated corresponding long-form answers with GPT-3.5 and Alpaca. We selected the datasets to cover characteristics of realistic RAG systems including specific prompt templates (present in RAGTruth, LFQA-Verification) and various domains (present in all datasets, specifically in SummEdits). Details and links to the datasets can be found in Table 4.
@@
Baselines. We use state-of-the-art baselines: AlignScore (Zha et al., 2023) (RoBERTa-based with pre- and postprocessing), MiniCheck (Tang et al., 2024), and Vectara-2.1222https://docs.vectara.com/docs (both T5-based). These complex state-of-the-art models are trained using custom datasets (Vectara, MiniCheck) or use postprocessing (AlignScore). As a teacher model to assign initial score, we use the best-performing model from the “complex” category, which (unlike LLMs) allow easy access to uncertainty scores and have acceptable performance. We employ optuna [^3] https://optuna.org/ as a principled way of choosing the remaining hyperparameters $\lambda_{u}^{\prime}$, $\lambda_{d}$, and the other teacher model used to estimate entailment probabilities for augmentations in Eqn. 1. We perform 50 trials per dataset and use the ROC score of a fine-tuned DeBERTaV2 model on the small validation dataset as selection objective. In case limited budget for hyperparameter tuning is available, we recomment setting $\lambda_{u}^{\prime}{=}\lambda_{d}\in\left[20,50\right]$ which led to stable performance. Auto-GDA is run for two iterations on RAGTruth and one iteration on the other datasets, generating synthetic datasets between 1.3×1.3$\times$1.3 × and 2×2$\times$2 × the original dataset size. We found no improvements through further increasing dataset size. We also compare against several common UDA methods, including robustness-based approaches. Specifically, we implement DAPT (Gururangan et al., 2020), SiFT (He et al., 2020), and Deep-CORAL (Sun & Saenko, 2016) for further pretraining of the DeBERTa-V2 model.
@@ 5.1 Synthetic Data for NLI Model Fine-Tuning
We present the main results obtained with Auto-GDA in Table 1. Our results show that Auto-GDA is highly effective and improves performance in ROC-AUC scores of all tested models on all datasets.
@@ 5.2 Ablation Investigations
Components of the algorithm. We add the components of our algorithm individually and show how they successively increase performance in Table 2. In all ablations, we keep dataset size and other parameters constant. The biggest gain is achieved by fine-tuning on data created through few-shot prompting. We subsequently add data augmentations without applying our selection criterion, but instead selecting few-shot and augmented samples randomly. We observe that this decreases data quality, highlighting that data augmentation is only beneficial together with our filtering criterion. When we do so and apply data augmentation with our filtering step (corresponding to full Auto-GDA), this increases performance overall with one exception on the LFQA-Verification dataset (note however that performance here is already above the labeled data, so selection based on target data may draw the results toward the labeled data scores as well). As an upper baseline we are interested in the hypothetical performance reachable by fine-tuning on human-labeled samples and include it in Table 2. Considering the difference between the no fine-tuning models and the models fine-tuned on human-labeled data as the domain adaptation gap, expressing our results relative to these baselines indicates that we manage to close an impressive 96% of this gap.
@@ 5.3 Inference Efficiency
Linking to our motivational Figure 1, we study the efficiency of our models in Table 3. We compute NLI scores for 50 random samples from the respective datasets. We observe models in three categories: The most efficient models (Vectara to BART-large) have medium performance on the RAG datasets used in this work indicated by their ROC. On the other hand, models using sophisticated post-processing (AlignScore) perform better, but require about 2.5 times more inference time than our most successful DeBERTa model. Finally, LLMs via APIs require about 10-fold inference time, but result in highest performance. When we compare models trained with our approach, we observe LLM-level performance at about 10% of the inference time. Although our primary focus lies on inference, we report generation times for Auto-GDA in Table 16 (Appendix) for completeness.
@@ 6 Disussion and Conclusion
In this work, we show that synthetic data can successfully tackle the domain generalization gap for NLI models. We present Auto-GDA, an automatic approach for synthetic sample generation and selection overcoming the need for tedious heuristic or manual filtering and augmentation selection. Our results show that we can obtain models that perform on par with most powerful LLMs while having around 90% less inference time using our method. By generating synthetic data, we can provide a more comprehensive and tailored representation, allowing for greater control over the desired features. Our results also confirm the common intuition that generalization is increasingly hard with smaller models (Bhargava et al., 2021). This highlights that domain adaptation is particularly important when low latency at inference time is required, whereas general purpose models can be preferable when quick inference is no hard requirement. Pretrained models as AlignScore (Zha et al., 2023) and MiniCheck (Tang et al., 2024) strike a good balance between inference time and performance when generation and fine-tuning is not possible or too costly.
Limitations. In our study we assume that the distribution of evidence samples including the retrieved documents is readily available. In many real-world applications, this may not be the case. To address this, techniques like passage clustering and summarization, as explored in Sarthi et al. (2024), could be employed on the knowledge base to cover the diversity of evidence passages as a surrogate of this distribution. In general, we advise practitioners to obtain the most representative corpus of the target domain for best results. Currently, domain adaptation requires a model for each individual domain. Future work is required to further study if it is possible to adapt efficient models for multiple domains without performance degradation. In addition, an assessment including the entire RAG system with a retrieval component using frameworks such as RAGChecker (Ru et al., 2024) may result in insightful in future work.
@@ B.1 Deriving our Objective from a Distribution Matching Criterion
Parametrizing densities. We need to insert a suitable and tractable parametrizations of $p_{\text{cov}}$ and $p_{\theta}$. We start by modeling their marginals. To model $p_{\text{cov}}$ we chose an efficiently tractable density $p_{\text{cov}}({\bm{x}})$ can be defined via the nearest target feature vector in $\text{claim}(\mathcal{D}_{t,{\bm{e}}})$ by444we now assume ${\bm{x}}\in\mathcal{X}$ to be in a metric space. This can be achieved using an encoder mapping the textual input to real vectors.
@@ C.1 Implementation Details for Augmentation Strategies
In this section we provide additional details regarding the data augmentation strategies that we deploy in this work. We fully commit to open-sourcing our code to reproduce the experiments upon acceptance.
@@ C.3 Auto-GDA Details and Hyperparameters
Base models. As a basis for finetuning we use huggingface checkpoints for DeBERTaV2777https://huggingface.co/tasksource/deberta-base-long-nli, BART-large [^8] https://huggingface.co/facebook/bart-large-mnli and FLAN-T5999https://huggingface.co/sjrhuschlee/flan-t5-base-mnli.
@@ C.5 Baseline NLI Models
For the complex NLI model baselines, we use Vectara HHEM-2.1101010https://huggingface.co/vectara/hallucination_evaluation_model. The model cannot be easily fine-tuned because it uses custom code. Additionally we use Alignscore-base with the checkpoint found in this repository [^11] https://github.com/yuh-zha/AlignScore with the recommended “split” (pre- and postprocessing) nli_sp option. We neglect the larger version as its runtime was comparable to LLMs at a usually lower performance, making the smaller model a better trade-off. Finally we use the best-performing Minicheck flan-t5-large model by Tang et al. (2024) from the official huggingface page [^12] https://huggingface.co/lytang/MiniCheck-Flan-T5-Large.
@@ C.6 Robust Pre-training and Fine-tuning for Unsupervised Domain Adaptation
Implementation Details For our robust optimization experiments we used the DeBERTaV2 based NLI model, and limited maximum tokenization length at 1,024 tokens across all benchmarks. For DAPT we extracted the DeBERTaV2 backbone and trained on the target domain for 1 full epoch, using 10% masking probability. At the fine-tuning stage of both methods we ran 1 full epoch on the MNLI dataset. At the masked pre-training and fine-tuning stages of the experiments, we used the AdamW optimizer with learning rate $10^{-5}$ and weight decay $10^{-3}$, enabling 100 warm-up steps over the supervised fine-tuning. For SiFT and CORAL we set the coefficient of the respective regularization terms to 0.5, after running hyperparameter optimization with a coarse grid search. Batch size used for covariance estimation in CORAL was set at 64. Each experiment was repeated over 3 random trials.
@@ D.3 Different teacher models
D.4 Robustness to Initial Data Quality
The initial data plays an essential role in Auto-GDA, however we designed our algorithm to be as fault-tolerant as possible and to be able to cope with some low-quality samples in the initial population. For instance, we do not solely rely on generative models, but use discriminative teacher models in the generation loop as well. The selection objective is specifically designed to filter out low-quality samples. Mislabeled samples in the fine-tuning dataset can severely affect performance of a fine-tuned model (Wang et al., 2023b). To test the robustness of Auto-GDA, we manually flip 50 % of the labels in the initial data to study its effect and trace these samples through the generation process. We use the LFQA-Verification dataset and perform 5 runs with different initial generation seeds (using the same hyperparameters as in Table 1 otherwise). First, while 50% of the data have flipped labels initially, in the data selected after one iteration of Auto-GDA, only 10.0% +- 1.1% are the initial data or augmentations of the data with the flipped labels. After fine-tuning on the generated data, the ROC-AUC score drops slightly by 1.1 points as shown in table below.
Original Performance with 50% flipped initial labels
ROC-AUC of fine-tuned DeBERTaV2 0.925 $\pm$ 0.009 0.914 $\pm$ 0.007
Despite the substantial amount of mislabeled data (50 %) the drop in performance remains small, highlighting that Auto-GDA is quite robust to data quality issues due to its fault-tolerant design.
D.5 Simple Baselines
@@
Dataset RAGTruth LFQA-Verif. SummEdits Mean
@@
Table 12: Fine-tuning on validation set does increase performance but does not reach Auto-GDAs performance on average.
D.6 Effect of Dataset Size
@@
We also study the effect of generating more LLM samples on the performance of the baseline of solely fine-tuning on LLM-generated data. We generate more LLM data using our few-shot prompting strategy and fine-tune the DeBERTa model on LLM-generated datasets in the range of 0.5x - 8x the size of the datasets used in the main paper. We obtain the results shown in Table 13 (ROC-Scores of the DeBERTaV2 model fine-tuned on the data). Perhaps surprisingly, increasing the number of generated samples does not increase the performance of fine-tuned models. We attribute this to the observation that LLMs (even when setting a higher temperature) will generate highly similar samples after a certain dataset size is reached, not further improving performance of a fine-tuned model. Performance even decreases again, potentially due to overfitting on the synthetic data. We therefore conclude that even with more LLM data, the performance does not reach the level of Auto-GDA.
dataset size 0.5x 1x 2x 4x 8x
LFQA 0.916 $\pm$ 0.007 0.924 $\pm$ 0.006 0.903 $\pm$ 0.013 0.896 $\pm$ 0.007 0.868 $\pm$ 0.003
SummEdits 0.880 $\pm$ 0.003 0.876 $\pm$ 0.002 0.876 $\pm$ 0.005 0.868 $\pm$ 0.002 0.855 $\pm$ 0.005
RAGTruth-Summary 0.821 $\pm$ 0.003 0.808 $\pm$ 0.005 0.791 $\pm$ 0.007 0.797 $\pm$ 0.001 0.783 $\pm$ 0.004
RAGTruth-QA 0.808 $\pm$ 0.007 0.828 $\pm$ 0.006 0.830 $\pm$ 0.006 0.834 $\pm$ 0.004 0.821 $\pm$ 0.005
average 0.860 0.856 0.849 0.844 0.841
Table 13: Results when fine-tuning on more LLM generated data (ROC-AUC scores for DeBERTaV2) show that performance of this baseline does not improve with more data.
Figure 8: Testing different synthetic dataset sizes and learning rates for fine-tuning. While the original dataset size is at about 3 claim examples per evidence (dashed line) we test a wide range of dataset sizes ranging from 1/50 examples per evidence (100x smaller than original) to larger datasets with 30 claims per evidence, (10x larger than original). We observe a relatively stable optimum from about a 1/3 of original size to twice the original size using learning rates around $10^{-5}$.
D.7 Ablating terms in the selection criterion
To better understand the effect of the different terms in our objective, we perform an additional ablation study only using a selection objective with one term at a time. We provide the results in Table 14. Despite some difference being small, we see that all terms contribute to the final performance.
dataset all terms only distance only label-cert. only utility
LFQA 0.925 $\pm$ 0.004 0.916 $\pm$ 0.004 0.919 $\pm$ 0.003 0.917 $\pm$ 0.005
Summedits 0.883 $\pm$ 0.002 0.877 $\pm$ 0.004 0.866 $\pm$ 0.002 0.881 $\pm$ 0.002
RAGTruth-Summary (1 run) 0.861 0.805 0.860 0.587
RAGTruth-QA (1 run) 0.827 0.756 0.814 0.823
Table 14: Ablation on terms in the selection criterion (ROC-AUC scores for DeBERTaV2). Using all three terms results in better performance than only using one of the terms each.
D.8 Effect of Source Data on UDA effectivenss